OpenStudy (anonymous):

1= (7/a-3) - (42/a^2-9)

OpenStudy (anonymous):

(7/a-3)-(42/a^2-9) here a^2-9=(a+3)(a-3) take L.C.M answer = (7a-21)/(a^2-9)

OpenStudy (anonymous):

what do you do with the 1 in front of the equation

OpenStudy (anonymous):

wait a mint

OpenStudy (anonymous):

ok write a^2-9=(a+3)(a-3) multiply both side of equation by (a+3)(a-3) at the equation becomes a^2-7a+21=0

OpenStudy (anonymous):

Oh ok! I gotcha, thank you so much!!